# references.bib

### v1

@@ 1 Introduction
Anomaly detection in time-series data has emerged as a common problem with a variety of real-world applications, from the medical field [moody_impact_2001] to high-performance computing [su_robust_2019]. Findings from the field of anomaly detection are also of particular interest to the modern automotive industry. This industry is very diverse, including, for example, manufacturing, prototyping and testing. Furthermore, a car is a complex structure that can be segmented into different subsystems, one of them being the powertrain, which includes all components required for longitudinal dynamics. The powertrain is an important subsystem, as it is something a customer interacts with when driving or being driven. Therefore, testing the powertrain is an an integral part of the wider automotive powertrain development and is undertaken at different stages of development. Each of these stages is composed of many integration levels. These integration levels range from powertrain sub-component testing, such as the electric drive unit (EDU) controller or high-voltage battery (HVB) management system, to whole vehicle powertrain testing. For each integration level there is a special type of controlled environment, called a test bench. The use-case in this paper is on an endurance powertrain test bench, where the EDU and HVB on their own are tested under different conditions and loads for longer periods to simulate wear over time. Given the costly maintenance and upkeep costs of such test benches, it is desirable to keep downtime at a minimum and to avoid faulty records. Also, it is desirable to detect problems early to prevent damage to the testee. Online time-series anomaly detection is especially relevant, as it can provide timely insights into potentially harming behaviour that deviates from the norm.
@@
To achieve this, we propose a temporal multi-head attention-based variational autoencoder (TeVAE). TeVAE consists of a bidirectional long short-term memory (BiLSTM) variational autoencoder architecture that maps a time-series window into a temporal latent distribution [park_multimodal_2018] [su_robust_2019]. Also, a multi-head attention (MA) mechanism is added to further enhance the sampled latent matrix before it is passed on to the decoder. As shown in the ablation study, this approach avoids the so-called bypassed phenomenon [bahuleyan_variational_2018], which is the first contribution. Furthermore, this paper offers a unique methodology for the reverse-window process. It is used for remapping the fixed-length windows the model is trained on to continuous variable-length sequences, i.e. time series. Moreover, we propose a set of metrics apt to online time-series anomaly detection that is not only interpretable and simple but is also compatible with discrete time-series anomaly detection. Lastly, the root-cause capability of TeVAE is investigated, for which a new metric, the root-cause precision, is also proposed.
@@ 2.2 Data Set
Table 1: Channels (features) chosen for modelling testee behaviour [correia_ma-vae_2023]. Indices correspond to Figure 1.
@@ 13 Inverter Temperature
Due to the absence of labelled anomalies in the dataset, realistic anomalous events are intentionally simulated and recorded following the advice of test bench engineers. To this end, four anomaly types are recorded. Every anomaly type is recorded for every drive cycle at least once, leading to $N_{\text{a}}=47$ anomalous measurements that are all used as the labelled anomalous portion of the testing subset $\mathcal{D}^{\text{test}}$. Hence $\mathcal{D}^{\text{test}}$ is made up of $N=745$ measurements, representing an anomaly ratio of around $N_{\text{a}}/N=6.3\%$, however in reality this value is estimated to be much lower. This amount of anomalous data in relation to anomaly-free data is used as it approximately matches the anomaly ratio in public data sets [mathur_swat_2016, ahmed_wadi_2017, hundman_detecting_2018, su_robust_2019] and because the data set is not large enough to create a larger anomaly-free test subset.
@@
Clearly this data set consists of several discrete time series. This is in contrast to continuous anomaly detection, which is defined as detecting anomalies in a process that exists for a longer continuous time period without breaks and is the most common type present in public data sets. This includes monitoring applications like in water distribution [mathur_swat_2016, ahmed_wadi_2017] or server machines [su_robust_2019]. Use cases of continuous anomaly detection tend to consist of a singular longer time series which contains anomaly-free and anomalous sub-sequences within it. On the other hand, we define discrete anomaly detection as detecting anomalies in chunks of processes that happen independently of each other, such as automotive test benches, where several tests may occur one after another but are not temporally contiguous and hence provide a multivariate time series for each test. Here the testees are not monitored over a longer period of time, but rather the measured time series are evaluated as they are being recorded. Arguably, analysis of heartbeat rhythms could be considered a discrete anomaly detection problem if applied generically to a number of test subjects, rather than individually, as the training and testing subsets each likely consist of multiple time series from different subjects. Therefore, data sets for discrete anomaly detection consist of several anomaly-free and anomalous time series, where a given anomalous time series may be entirely anomalous or only partly.
@@
Given that some channels (such as torque) are sampled much faster than others (like temperature and SoC), a common sampling rate of $2$Hz is chosen. Channels sampled slower than $2$Hz are linearly interpolated, which is considered permissible due to the lower amplitude resolution of those channels. Channels sampled faster than $2$Hz are passed through a low-pass filter with a cut-off frequency of $1$Hz and then resampled to $2$Hz, as is consistent with the Whittaker–Nyquist–Shannon theorem [shannon_communication_1949]. Then, the measurements are z-score normalised, i.e. transformed such that the mean for each channel lies at 00 and the standard deviation at $1$. Lastly, the measurements are windowed to create a set of fixed-length sub-sequences, or windows. The reasoning behind using windows rather than entire sequences is that most of the dynamic processes in the time-series data are fairly fast and hence only have short-term effect that lasts for a small period of time. Modelling entire variable-length sequences would be possible but much of the model learning capacity would be wasted on internal state mechanisms carrying information for longer periods of time than necessary. This therefore allows for more efficient model training on windows that are only as long as they need to be to represent all present dynamics in the time-series data. To find the effect length of the slowest dynamics an autocorrelation analysis is undertaken for each of the drive cycles and each of the features within them. This is in contrast to approaches taken in literature, which treat window size as a hyperparameter that is tuned. It is unclear how hyperparameter tuning can be undertaken outside of a supervised problem, hence it is assumed that in the real world, this process would not be possible. An autocorrelation analysis yields the correlation between the time series and a range of lagged versions of the time series for a specified number of lags and does not require labels. An example of an autocorrelation plot for an arbitrary measurement and arbitrary channel is shown in Figure 2.
@@
In an operative environment, it is desirable to find out whether the previously recorded sequence had any problems to analyse before the next measurement is recorded. Also, a model that performs as well as possible with as little data as possible translates to faster deployment. Good performance is indicated by a model that can detect as many anomalies as possible and rarely labels anomaly-free measurements wrongly. To investigate the required training subset size of the model, it is trained with $1$h, $8$h, $64$h, and $512$h of dynamic testing time, which corresponds to the first $6$, $44$, $348$, and $2785$ measurements, respectively. The results are also represented in Section 5. In each of the above-mentioned cases, the unlabelled training subset $\mathcal{D}^{\text{train}}$ is further split into a training ($80\%$) and a validation ($20\%$) subsets.
@@ 2.3 Variational Autoencoders
The variational autoencoder [kingma_auto-encoding_2014][rezende_stochastic_2014] is a generative model that structurally resembles an autoencoder, but is theoretically derived from variational Bayesian statistics. As opposed to the regular deterministic autoencoder, the VAE uses the evidence lower bound (ELBO), which is a lower bound approximation of the so-called log evidence $\log p_{\theta}(\textbf{X})$, as its objective function. The ELBO, Equation 3, can be expressed as the reconstruction log-likelihood and the negative Kullback-Leibler Divergence ($D_{\text{KL}}$) between the approximate posterior $q_{\phi}(\textbf{Z}|\textbf{X})$ and the prior $p_{\theta}(\textbf{Z})$, which is typically assumed to be a Gaussian distribution [goodfellow_deep_2016]. The training data does not have to follow a Gaussian distribution since the latent distribution is a nonlinear mapping of the training data.
$\begin{split}\mathcal{L}_{\theta,\phi}(\textbf{X})&=\mathbb{E}_{\textbf{Z}\sim q_{\phi}(\textbf{Z}|\textbf{X})}\left[\log p_{\theta}(\textbf{X}|\textbf{Z})\right]\\ &-D_{\text{KL}}(q_{\phi}(\textbf{Z}|\textbf{X})||p_{\theta}(\textbf{Z}))\\ \end{split}$ (3)
@@
Finally, to enable the backpropagation through the otherwise intractable gradient of the ELBO, the reparametrisation trick [kingma_auto-encoding_2014] is applied, shown in Equation 5. This is one of the reasons for the use of a Gaussian distribution as the latent distribution, although any distribution type from the "location-scale" type can be used [kingma_auto-encoding_2014], like a Laplace distribution.
@@ 2.4 Multi-head Attention Mechanism
MS consists of two different concepts: self-attention and its multi-head extension. Self-attention is nothing more than scaled dot-product attention [vaswani_attention_2017] where the key, query and value are the same. The scaled dot-product attention score is the softmax [bridle_probabilistic_1990] of the product between query matrix Q and key matrix K which is scaled by $\sqrt{d_{\textbf{K}}}$. The product between the attention score and the value matrix V yields the context matrix C, as shown in Equation 6.
@@
Compared to recurrent or convolutional layers, self-attention offers a variety of benefits, such as the reduction of computational complexity, as well as an increased amount of operations that can be parallelised [vaswani_attention_2017]. Also, self-attention inherits an advantage over Bahdanau-style attention [bahdanau_neural_2015] from the underlying scaled dot-product attention mechanism: it can run efficiently in matrix multiplication manner [vaswani_attention_2017].
Multi-head self-attention then allows the attention model to attend to different representation subspaces [vaswani_attention_2017], in addition to learning useful projections rather than it being a stateless transformation [chollet_deep_2021]. This is achieved using weight matrices $\textbf{W}^{Q}_{i}$, $\textbf{W}^{K}_{i}$, $\textbf{W}^{V}_{i}$, which contain trainable parameters and are unique for each head $i$, as shown in Equation 7.
@@
The underlying mechanism of MA is identical to MS, with the only difference being that K $=$ Q $\not$= V. Essentially, MA finds which time steps correlate most with each other inside a given input window and weighs the time steps in context matrix C accordingly. The benefit of this alteration is discussed in Section 4.
@@ 3 Related Work
In time-series anomaly detection literature, the only other model that uses the combination of a VAE and an attention mechanism is by [pereira_unsupervised_2018]. For the purpose of our paper, it is referred to as variational self-attention VAE (VS-VAE). Their approach consists of a BiLSTM encoder and decoder, where, for an input window of length $w$, the $t=w$ encoder hidden states of each direction are passed on to the variational self-attention (VS) mechanism [bahuleyan_variational_2018]. The resulting context vector is then concatenated with the sampled latent vector and then passed on to the decoder. The author claims that applying VS to the VAE model solves the bypass phenomenon, however, no evidence for this claim is provided.
The first published time-series anomaly detection approach based on VAE is LSTM-VAE [park_multimodal_2018]. One of the contributions is its use of a dynamic prior, i.e. $\mathcal{N}(\mu_{p},1)$, rather than a static one, i.e. $\mathcal{N}(0,1)$. In addition to that, they introduce a state-based threshold estimation method consisting of a support-vector regressor (SVR), which maps the latent distribution parameters $(\mu_{\textbf{z}},\sigma_{\textbf{z}})$ to the resulting anomaly score using the validation data. Hence, the dynamic threshold can be obtained through Equation 10.
@@
OmniAnomaly [su_robust_2019] attempts to create a temporal connection between latent distributions by applying a linear Gaussian state space model to them. For the purpose of this paper, it is abbreviated to OmniA. Also, it concatenates the last gated recurrent unit (GRU) hidden state with the latent vector sampled in the previous time step. In addition to that, it uses planar normalising flow [rezende_variational_2015] by applying $K$ transformations to the latent vector in order to approximate a non-Gaussian posterior, as shown in Equation 11.
@@
A simplified VAE architecture [pereira_unsupervised_2019] based on BiLSTM layers is also proposed. For the purpose of our paper, it is called Wasserstein VAE (W-VAE). Unlike its predecessor [pereira_unsupervised_2018], it drops the attention mechanism but provides contributions elsewhere. It offers two strategies to detect anomalies based on the VAE outputs. The first involves clustering the space characterised by the mean parameter of the latent distribution into two clusters and labelling the larger one as normal. This strategy has a few weaknesses: it cannot be used in an operative environment as it requires some sort of history of test windows to form the clusters and it assumes that there are always anomalous samples present. The second strategy finds the Wasserstein similarity measure between the latent mean space mapping of the test window in question and the respective mapping $i$ resulting from a representative data subset, such as the validation subset. Equation 12 shows how the Wasserstein similarity measure is computed
@@
Sliding-window convolutional variational autoencoder (SWCVAE) [chen_unsupervised_2020] is the first that applies convolutional neural networks (CNN) to VAEs for multivariate time-series anomaly detection. Peculiarly, 2D CNN layers are used with the justification of being able to process the input both spatially and temporally. We, however, doubt the ability of the model to properly detect anomalies through spatial processing, as a kernel moving along the feature axis can only capture features adjacent to each other. To create a continuous anomaly score from windows they append the last value of each window to the previous one. For the purpose of this paper, this process is referred to as last-type reverse-windowing.
Smoothness-inducing sequential VAE (SISVAE) [li_anomaly_2021] tries to improve the modelling robustness by the addition of a smoothing term in the loss function which contributes to the reduction of sudden changes in the reconstructed signal, making it less sensitive to noisy time steps.
As part of the variational autoencoder-based selective prediction (VASP) framework [von_schleinitz_vasp_2021], a variational autoencoder architecture is proposed to increase the robustness of time-series prediction when faced with anomalies. While the main contribution is attributed to the framework itself, not the VAE, it should be noted that during inference only the mean parameter of the latent distribution is passed to the decoder.
FedAnomaly [zhang_federated_2021] first extends time-series anomaly detection to a federated learning setting, which may be relevant in use cases involving large amounts of data or strict privacy settings. Other than being the first VAE based on convolutional GRU layers, the contributions of this paper lie exclusively within the federated framework. Results show that as a single entity, it performs better than in a federated scenario.
Other than a feature selection procedure based on the Kolmogorow-Smirnow test, the lightweight LSTM-VAE (LW-VAE) [fahrmann_lightweight_2022] offers no significant contributions besides being relatively parameter-light.

### v2

# TeVAE: A Variational Autoencoder Approach for Discrete Online Anomaly Detection in Variable-state Multivariate Time-series Data

@@ 1 Introduction
Anomaly detection in time-series data has emerged as a common problem with a variety of real-world applications, from the medical field [16] to high-performance computing [24]. Findings from the field of anomaly detection are also of particular interest to the modern automotive industry. This industry is very diverse, including, for example, manufacturing, prototyping and testing. Furthermore, a car is a complex structure that can be segmented into different subsystems, one of them being the powertrain, which includes all components required for longitudinal dynamics. The powertrain is an important subsystem, as it is something a customer interacts with when driving or being driven. Therefore, testing the powertrain is an an integral part of the wider automotive powertrain development and is undertaken at different stages of development. Each of these stages is composed of many integration levels. These integration levels range from powertrain sub-component testing, such as the electric drive unit (EDU) controller or high-voltage battery (HVB) management system, to whole vehicle powertrain testing. For each integration level there is a special type of controlled environment, called a test bench. The use-case in this paper is on an endurance powertrain test bench, where the EDU and HVB on their own are tested under different conditions and loads for longer periods to simulate wear over time. Given the costly maintenance and upkeep costs of such test benches, it is desirable to keep downtime at a minimum and to avoid faulty records. Also, it is desirable to detect problems early to prevent damage to the testee. Online time-series anomaly detection is especially relevant, as it can provide timely insights into potentially harming behaviour that deviates from the norm.
@@
To achieve this, we propose a temporal multi-head attention-based variational autoencoder (TeVAE). TeVAE consists of a bidirectional long short-term memory (BiLSTM) variational autoencoder architecture that maps a time-series window into a temporal latent distribution [17] [24]. Also, a multi-head attention (MA) mechanism is added to further enhance the sampled latent matrix before it is passed on to the decoder. As shown in the ablation study, this approach avoids the so-called bypassed phenomenon [3], which is the first contribution. Furthermore, this paper offers a unique methodology for the reverse-window process. It is used for remapping the fixed-length windows the model is trained on to continuous variable-length sequences, i.e. time series. Moreover, we propose a set of metrics apt to online time-series anomaly detection that is not only interpretable and simple but is also compatible with discrete time-series anomaly detection. Lastly, the root-cause capability of TeVAE is investigated, for which a new metric, the root-cause precision, is also proposed.
@@ 2.2 Data Set
Table 1: Channels (features) chosen for modelling testee behaviour [7]. Indices correspond to Figure 1.
@@ 13 Inverter Temperature
Due to the absence of labelled anomalies in the dataset, realistic anomalous events are intentionally simulated and recorded following the advice of test bench engineers. To this end, four anomaly types are recorded. Every anomaly type is recorded for every drive cycle at least once, leading to $N_{\text{a}}=47$ anomalous measurements that are all used as the labelled anomalous portion of the testing subset $\mathcal{D}^{\text{test}}$. Hence $\mathcal{D}^{\text{test}}$ is made up of $N=745$ measurements, representing an anomaly ratio of around $N_{\text{a}}/N=6.3\%$, however in reality this value is estimated to be much lower. This amount of anomalous data in relation to anomaly-free data is used as it approximately matches the anomaly ratio in public data sets [15, 1, 11, 24] and because the data set is not large enough to create a larger anomaly-free test subset.
@@
Clearly this data set consists of several discrete time series. This is in contrast to continuous anomaly detection, which is defined as detecting anomalies in a process that exists for a longer continuous time period without breaks and is the most common type present in public data sets. This includes monitoring applications like in water distribution [15, 1] or server machines [24]. Use cases of continuous anomaly detection tend to consist of a singular longer time series which contains anomaly-free and anomalous sub-sequences within it. On the other hand, we define discrete anomaly detection as detecting anomalies in chunks of processes that happen independently of each other, such as automotive test benches, where several tests may occur one after another but are not temporally contiguous and hence provide a multivariate time series for each test. Here the testees are not monitored over a longer period of time, but rather the measured time series are evaluated as they are being recorded. Arguably, analysis of heartbeat rhythms could be considered a discrete anomaly detection problem if applied generically to a number of test subjects, rather than individually, as the training and testing subsets each likely consist of multiple time series from different subjects. Therefore, data sets for discrete anomaly detection consist of several anomaly-free and anomalous time series, where a given anomalous time series may be entirely anomalous or only partly.
@@
Given that some channels (such as torque) are sampled much faster than others (like temperature and SoC), a common sampling rate of 22Hz is chosen. Channels sampled slower than 22Hz are linearly interpolated, which is considered permissible due to the lower amplitude resolution of those channels. Channels sampled faster than 22Hz are passed through a low-pass filter with a cut-off frequency of 11Hz and then resampled to 22Hz, as is consistent with the Whittaker–Nyquist–Shannon theorem [23]. Then, the measurements are z-score normalised, i.e. transformed such that the mean for each channel lies at 0 and the standard deviation at $1$. Lastly, the measurements are windowed to create a set of fixed-length sub-sequences, or windows. The reasoning behind using windows rather than entire sequences is that most of the dynamic processes in the time-series data are fairly fast and hence only have short-term effect that lasts for a small period of time. Modelling entire variable-length sequences would be possible but much of the model learning capacity would be wasted on internal state mechanisms carrying information for longer periods of time than necessary. This therefore allows for more efficient model training on windows that are only as long as they need to be to represent all present dynamics in the time-series data. To find the effect length of the slowest dynamics an autocorrelation analysis is undertaken for each of the drive cycles and each of the features within them. This is in contrast to approaches taken in literature, which treat window size as a hyperparameter that is tuned. It is unclear how hyperparameter tuning can be undertaken outside of a supervised problem, hence it is assumed that in the real world, this process would not be possible. An autocorrelation analysis yields the correlation between the time series and a range of lagged versions of the time series for a specified number of lags and does not require labels. An example of an autocorrelation plot for an arbitrary measurement and arbitrary channel is shown in Figure 2.
@@
In an operative environment, it is desirable to find out whether the previously recorded sequence had any problems to analyse before the next measurement is recorded. Also, a model that performs as well as possible with as little data as possible translates to faster deployment. Good performance is indicated by a model that can detect as many anomalies as possible and rarely labels anomaly-free measurements wrongly. To investigate the required training subset size of the model, it is trained with 11h, 88h, 6464h, and 512512h of dynamic testing time, which corresponds to the first $6$, $44$, $348$, and $2785$ measurements, respectively. The results are also represented in Section 5. In each of the above-mentioned cases, the unlabelled training subset $\mathcal{D}^{\text{train}}$ is further split into a training ($80\%$) and a validation ($20\%$) subsets.
@@ 2.3 Variational Autoencoders
The variational autoencoder [12][21] is a generative model that structurally resembles an autoencoder, but is theoretically derived from variational Bayesian statistics. As opposed to the regular deterministic autoencoder, the VAE uses the evidence lower bound (ELBO), which is a lower bound approximation of the so-called log evidence $\log p_{\theta}(\textbf{X})$, as its objective function. The ELBO, Equation 3, can be expressed as the reconstruction log-likelihood and the negative Kullback-Leibler Divergence ($D_{\text{KL}}$) between the approximate posterior $q_{\phi}(\textbf{Z}|\textbf{X})$ and the prior $p_{\theta}(\textbf{Z})$, which is typically assumed to be a Gaussian distribution [10]. The training data does not have to follow a Gaussian distribution since the latent distribution is a nonlinear mapping of the training data.
[…]
@@
Finally, to enable the backpropagation through the otherwise intractable gradient of the ELBO, the reparametrisation trick [12] is applied, shown in Equation 5. This is one of the reasons for the use of a Gaussian distribution as the latent distribution, although any distribution type from the "location-scale" type can be used [12], like a Laplace distribution.
@@ 2.4 Multi-head Attention Mechanism
MS consists of two different concepts: self-attention and its multi-head extension. Self-attention is nothing more than scaled dot-product attention [25] where the key, query and value are the same. The scaled dot-product attention score is the softmax [4] of the product between query matrix Q and key matrix K which is scaled by $\sqrt{d_{\textbf{K}}}$. The product between the attention score and the value matrix V yields the context matrix C, as shown in Equation 6.
@@
Compared to recurrent or convolutional layers, self-attention offers a variety of benefits, such as the reduction of computational complexity, as well as an increased amount of operations that can be parallelised [25]. Also, self-attention inherits an advantage over Bahdanau-style attention [2] from the underlying scaled dot-product attention mechanism: it can run efficiently in matrix multiplication manner [25].
Multi-head self-attention then allows the attention model to attend to different representation subspaces [25], in addition to learning useful projections rather than it being a stateless transformation [6]. This is achieved using weight matrices $\textbf{W}^{Q}_{i}$, $\textbf{W}^{K}_{i}$, $\textbf{W}^{V}_{i}$, which contain trainable parameters and are unique for each head $i$, as shown in Equation 7.
@@
The underlying mechanism of MA is identical to MS, with the only difference being that K == Q $\not$= V. Essentially, MA finds which time steps correlate most with each other inside a given input window and weighs the time steps in context matrix C accordingly. The benefit of this alteration is discussed in Section 4.
@@ 3 Related Work
In time-series anomaly detection literature, the only other model that uses the combination of a VAE and an attention mechanism is by [18]. For the purpose of our paper, it is referred to as variational self-attention VAE (VS-VAE). Their approach consists of a BiLSTM encoder and decoder, where, for an input window of length $w$, the $t=w$ encoder hidden states of each direction are passed on to the variational self-attention (VS) mechanism [3]. The resulting context vector is then concatenated with the sampled latent vector and then passed on to the decoder. The author claims that applying VS to the VAE model solves the bypass phenomenon, however, no evidence for this claim is provided.
The first published time-series anomaly detection approach based on VAE is LSTM-VAE [17]. One of the contributions is its use of a dynamic prior, i.e. $\mathcal{N}(\mu_{p},1)$, rather than a static one, i.e. $\mathcal{N}(0,1)$. In addition to that, they introduce a state-based threshold estimation method consisting of a support-vector regressor (SVR), which maps the latent distribution parameters $(\mu_{\textbf{z}},\sigma_{\textbf{z}})$ to the resulting anomaly score using the validation data. Hence, the dynamic threshold can be obtained through Equation 10.
@@
OmniAnomaly [24] attempts to create a temporal connection between latent distributions by applying a linear Gaussian state space model to them. For the purpose of this paper, it is abbreviated to OmniA. Also, it concatenates the last gated recurrent unit (GRU) hidden state with the latent vector sampled in the previous time step. In addition to that, it uses planar normalising flow [20] by applying $K$ transformations to the latent vector in order to approximate a non-Gaussian posterior, as shown in Equation 11.
@@
A simplified VAE architecture [19] based on BiLSTM layers is also proposed. For the purpose of our paper, it is called Wasserstein VAE (W-VAE). Unlike its predecessor [18], it drops the attention mechanism but provides contributions elsewhere. It offers two strategies to detect anomalies based on the VAE outputs. The first involves clustering the space characterised by the mean parameter of the latent distribution into two clusters and labelling the larger one as normal. This strategy has a few weaknesses: it cannot be used in an operative environment as it requires some sort of history of test windows to form the clusters and it assumes that there are always anomalous samples present. The second strategy finds the Wasserstein similarity measure between the latent mean space mapping of the test window in question and the respective mapping $i$ resulting from a representative data subset, such as the validation subset. Equation 12 shows how the Wasserstein similarity measure is computed
@@
Sliding-window convolutional variational autoencoder (SWCVAE) [5] is the first that applies convolutional neural networks (CNN) to VAEs for multivariate time-series anomaly detection. Peculiarly, 2D CNN layers are used with the justification of being able to process the input both spatially and temporally. We, however, doubt the ability of the model to properly detect anomalies through spatial processing, as a kernel moving along the feature axis can only capture features adjacent to each other. To create a continuous anomaly score from windows they append the last value of each window to the previous one. For the purpose of this paper, this process is referred to as last-type reverse-windowing.
Smoothness-inducing sequential VAE (SISVAE) [14] tries to improve the modelling robustness by the addition of a smoothing term in the loss function which contributes to the reduction of sudden changes in the reconstructed signal, making it less sensitive to noisy time steps.
As part of the variational autoencoder-based selective prediction (VASP) framework [22], a variational autoencoder architecture is proposed to increase the robustness of time-series prediction when faced with anomalies. While the main contribution is attributed to the framework itself, not the VAE, it should be noted that during inference only the mean parameter of the latent distribution is passed to the decoder.
FedAnomaly [29] first extends time-series anomaly detection to a federated learning setting, which may be relevant in use cases involving large amounts of data or strict privacy settings. Other than being the first VAE based on convolutional GRU layers, the contributions of this paper lie exclusively within the federated framework. Results show that as a single entity, it performs better than in a federated scenario.
Other than a feature selection procedure based on the Kolmogorow-Smirnow test, the lightweight LSTM-VAE (LW-VAE) [8] offers no significant contributions besides being relatively parameter-light.